OpenStudy (phagey):

1. A vehicle is purchased for $19,800, and the lender requires a down payment of 20% of the purchase price. Find the down payment. $15,840 $7,023 $3,960 $396 2. The amount financed is the difference between the purchase price and the down payment. True False 3. A compact car is purchased for $12,800 with a down payment of $3,085. The balance in financed for three years at an annual rate of 5%. Calculate the monthly car payment. $92.46 $291.17 $383.63 $489.12

OpenStudy (tkhunny):

Show your work, please. Surely, you can do the first.

OpenStudy (phagey):

ok is number 1 $3,960 i multiplied 19,800 with 20

OpenStudy (tkhunny):

19800 * 20% = 19800 * 0.20 = 3960 -- Yup. Next.

OpenStudy (phagey):

2 is true

OpenStudy (phagey):

i dont know how to do number 3

OpenStudy (tkhunny):

Good. First, find the amount financed.

OpenStudy (phagey):

will i do 12,800-6400? to find the amount finaced

OpenStudy (tkhunny):

?? 12800 - 3085 = ??

OpenStudy (phagey):

how you get 3085?

OpenStudy (phagey):

but its 9715?

OpenStudy (tkhunny):

Ummm... The problem statement? A compact car is purchased for $12,800 with a down payment of $3,085. The balance in financed for three years at an annual rate of 5%. Calculate the monthly car payment. 152800 - 3085 = 9715

OpenStudy (phagey):

oh i got mixed up sorry whats our next step?

OpenStudy (tkhunny):

Either: 1) You have the right formula for periodic payments, or 2) We create the formula, or 3) Maybe we can use a spreadsheet. Take your pick.